OpenStudy (anonymous):

Frank plants a rectangular vegetable garden with a perimeter of 39 meters. The length is twice the width. Find the dimensions of the garden?

OpenStudy (anonymous):

39 * (39*2)

OpenStudy (lgbasallote):

let width = w length = 2w (since it's twice the width) the perimeter of the rectangle is 39 so \[39 = l + l + w + w \implies 39 = 2w + 2w + w + w \implies 39 = 6w\] solve for w

OpenStudy (anonymous):

w=6.5 ?

OpenStudy (lgbasallote):

seems so

OpenStudy (lgbasallote):

now solve for l